OpenStudy (megannicole51):

calculate the integral e^x/(e^x-2)(e^x+5)

OpenStudy (megannicole51):

@pgpilot326

OpenStudy (megannicole51):

@satellite73

OpenStudy (anonymous):

you are going to give me a complex first a u sub then partial fractions

OpenStudy (anonymous):

i.e. first compute the integral of \[\frac{x}{(x-2)(x+5)}\] then replace \(x\) by \(e^x\)

OpenStudy (megannicole51):

hahahah im sorry!!

OpenStudy (anonymous):

i refuse to solve using linear systems (just kidding)

OpenStudy (megannicole51):

wait you can replace e^x with x??

OpenStudy (megannicole51):

lol u know u love helping me!

OpenStudy (anonymous):

ok if you don't like that, put \(u=e^x, du =e^xdx\) and get \[\frac{u}{(u-2)(u+5)}\] if that is better on the eyes

OpenStudy (anonymous):

i just replaced \(e^x\) by \(x\) instead of by \(u\) but no matter

OpenStudy (megannicole51):

i like x better:)

OpenStudy (anonymous):

i bet you can finish from here right? it is going to be another partial fraction problem, the only difference was the u - sub as a first step

OpenStudy (megannicole51):

so when do i put back the e^x?

OpenStudy (anonymous):

\[\frac{x}{(x-2)(x+5)}=\frac{A}{x-2}+\frac{B}{x+5}\] at the very end, as any u - sub right?

OpenStudy (megannicole51):

right!

OpenStudy (megannicole51):

if i get it wrong then ill be back!!

OpenStudy (anonymous):

u - sub is the first technique you learned i am sure

OpenStudy (anonymous):

yeah but really really try to find A and B in your head i am sure you can do it, even though they are fractions

OpenStudy (megannicole51):

okay ill try!

OpenStudy (anonymous):

try it now and let me know what you get

OpenStudy (anonymous):

you two...

OpenStudy (megannicole51):

okay i got A=(2/7) and B=(5/7) and its wrong....what did i do:(

OpenStudy (megannicole51):

jk i know whats wrong!!! i didnt substitute it back!!

OpenStudy (megannicole51):

nope its still wrong:( lol

OpenStudy (anonymous):

no i think that is right

OpenStudy (anonymous):

oh no i am so sorry, i steered you wrong

OpenStudy (anonymous):

it is not \[\frac{x}{(x-2)(x+5)}\] it is \[\frac{1}{(x-2)(x+5)}\] damn damn damn

OpenStudy (megannicole51):

(2/7)ln(abs(e^x-2))+(5/7)ln(abs(e^x+5))+c

OpenStudy (anonymous):

many apologies, it must be getting late i did the damn u sub wrong!!

OpenStudy (anonymous):

\[u=e^x, du =e^xdx\] gives \[\frac{1}{(x-2)(x+5)}\] not matter, we can fix it without any more work

OpenStudy (anonymous):

just erase the 2 and the 5 in the numerators and replace them by 1

OpenStudy (anonymous):

(1/7)ln(abs(e^x-2))+(1/7)ln(abs(e^x+5))+c

OpenStudy (anonymous):

i am really sorry. all that work for nothing

OpenStudy (megannicole51):

its still wrong:/

OpenStudy (megannicole51):

no dont be sorry! im getting practice doing the steps!

OpenStudy (anonymous):

ok hold the phone

OpenStudy (anonymous):

\[\frac{1}{(x-2)(x+5)}=\frac{A}{x-2}+\frac{B}{x+5}\] put \(x=2\) get \(A+\frac{1}{7}\)

OpenStudy (anonymous):

mean get \(A=\frac{1}{7}\)

OpenStudy (anonymous):

put \(x=-5\) get \(B=-\frac{1}{7}\)

OpenStudy (anonymous):

try \[\frac{1}{7}\ln(e^x-2)-\frac{1}{7}\ln(e^x+5)\]

OpenStudy (anonymous):

plus C and absolute values etc

OpenStudy (megannicole51):

yay!!!!

OpenStudy (megannicole51):

thank you thank you!!

OpenStudy (anonymous):

whew sorry btw did you see how easy it is to find A and B? i mean the second time...

OpenStudy (megannicole51):

yeah i did! i just need to practice more

OpenStudy (anonymous):

good, do it, it is easy eventually